3 added to 5times a number and the result multiplied by 4 the final result72
Answer by addingup(3677)

multiplied by 4= 72. Do it backward, 72/4= 18
3 added to 5= 8
You cannot get to 18 from 8 (8*2=16; 8*3= 24)
Please check the numbers and re-post